OpenStudy (anonymous):

Help me please 4 questions (Pre-Algebra and Exponents)!!! and thank you ^_^

OpenStudy (anonymous):

Simplify 6 to the 5th over 7 to the 3rd all raised to the 2nd power. 6^7 over 7^10 6^10 over 7^6 6^3 over 7 12^5 over 14^3

OpenStudy (anonymous):

Do I have to add or what to the exponents?

OpenStudy (solomonzelman):

\(\large\color{black}{ \left(\begin{matrix}\frac{\LARGE 6^5}{\LARGE 7^3} \\ \end{matrix}\right)^2 }\)

OpenStudy (solomonzelman):

like this?

OpenStudy (anonymous):

correct

OpenStudy (solomonzelman):

you do not add or subtract exponents, because you have 7 on the bottom and 6 on the top. They are difference number. you would have been subtracting exponents if you have 6 on the bottom instead of 7.

OpenStudy (solomonzelman):

We do know that: \(\large\color{black}{ \left(\begin{matrix}\frac{\LARGE 6^5}{\LARGE 7^3} \\ \end{matrix}\right)^2 ~~~\Rightarrow ~~~\frac{\LARGE (6^5)^2}{\LARGE (7^3)^2} }\) and use a rule: \(\large\color{black}{ (a^b)^{~c}=a^{b ~\times~ c} }\) for top and bottom.

OpenStudy (anonymous):

Ok Ill try it!! :D

OpenStudy (solomonzelman):

go ahead. and take your time if you need to.

OpenStudy (solomonzelman):

the connection is bad for me too.

OpenStudy (solomonzelman):

(I am here, so many glitches) anyway, what do you get? (if you need help, ask....)

OpenStudy (anonymous):

I got...